### v1

# Atomic altermagnetism

Abstract
Altermagnetism has been recently experimentally verified by photoemission mapping of the spin order in momentum space in MnTe and CrSb, which feature two anisotropic sublattices with antiparallel magnetic dipole moments. In this work, we explicitly demonstrate the presence of an even-parity ferroically ordered non-dipolar spin density on atomic sites, i.e. atomic altermagnetism, in MnTe, La2O3Mn2Se2 and Ba2CaOsO6. We do so through spin-symmetry analysis and partial-wave decomposition of the spin density obtained by first-principles calculations. In MnTe we show a ferroically ordered $g$-wave form factor in the spin density around the Mn site. In the A2O3M2Se2 family (A= La, Sr and M= Mn, Fe, Co), we show that there is a ferroically ordered $d$-wave form factor coexisting with the antiferroic magnetic dipoles in the M site, while the O site shows no dipole but a pure $d$-wave atomic spin density. In the Mott-insulating Ba2CaOsO6, as a key result, we reveal a pure form of atomic altermagnetism – absent of any dipolar sublattice order. This highlights that the altermagnetic order can exist without a Néel vector formed by antiferroic dipole moments on an even number of crystal sublattices, underlining its distinction from collinear Néel antiferromagnetic order. Our calculations predict that La2O3Mn2Se2 and Ba2CaOsO6 can exhibit giant spin-splitter angles of up to 42° and 26° respectively, thus demonstrating the possibility of large altermagnetic responses without requiring the staggered Néel order of local dipole moments.
Figure 1: Partial-wave expansion of the spin density from non-relativistic DFT calculations on CrSb, A2O3M2Se2, and Ba2CaOsO6. (d), (e) and (f) show the radial dependence of the relevant partial-wave components $S_{lm}(r)$ as defined in the main text. (a) CrSb unit cell, where black arrows represent the magnetic dipoles and, superimposed, we show the ferroically ordered $g_{\bar{3}}=Y_{4,-3}$ harmonic. (d) Radial dependence of the order parameters. (top) Antiferroic $s$-wave on each sublattice, (bottom) ferroic $g$-wave form factor.(b) Unit cell of La2O3Mn2Se2 with an iso-surface of the spin density. Mn sites have opposite magnetic dipole moments, and O site has no dipole but a $d_{x^{2}-y^{2}}$ spin density. (e) (top) $s$-wave harmonic expansion centered at each Mn site. (bottom) $d$-wave direct-space order parameter, formed by adding up the $d_{x^{2}-y^{2}}$ contributions of MnA, MnB and O sublattices. (c) Total spin density iso-surface around the Os site in Ba2CaOsO6 within its octahedral environment. (f) shows the respective relevant partial-waves, with zero $s$-wave contribution (no magnetic dipole) and two ferroic $d$-wave harmonics $d_{xy}$ and $g_{\bar{2}}\equiv Y_{4,-2}$.
Altermagnetism is a recently identified new type of magnetism characterized by spin order in both direct and momentum space with $d$-, $g$-, or $i$-wave symmetry Šmejkal et al. (2022a). Altermagnetism was predicted by a rigorous classification and delimitation of collinear magnetic phases based on spin symmetries that involve pairs of generally distinct operations, acting on the lattice and spin degrees of freedom Šmejkal et al. (2022a). These spin symmetries lead to an unconventional complex altermagnetic spin density Šmejkal et al. (2020) and exchange fields Šmejkal et al. (2023), which break the underlying lattice symmetry in an analogous way to what occurs in unconventional superfluid states Šmejkal et al. (2022b); Jungwirth et al. (2024). Recently, altermagnetism has been experimentally confirmed by momentum-space and direct-space photoemission spectroscopy and microscopy in representative binary-compounds MnTe and CrSb with ordering above room temperature Krempaskỳ et al. (2024); Lee et al. (2024); Osumi et al. (2024); Reimers et al. (2024); Yang et al. (2024); Ding et al. (2024); Zeng et al. (2024); Li et al. (2024); Hariki et al. (2024); Amin et al. (2024).
The discovery of the altermagnetic spin symmetry class also provided a unifying explanation of the previous reports of unconventional time-reversal symmetry breaking (TRSB) in the electronic structure and related effects, such as the anomalous Hall effect Šmejkal et al. (2020); Mazin et al. (2021), spin currents Ahn et al. (2019); Naka et al. (2019); González-Hernández et al. (2021); Šmejkal et al. (2022c) and magneto-optical effects Samanta et al. (2020); Mazin et al. (2021). The unconventional electronic structure of altermagnets exhibits TRSB spin-split bands across the entire Brillouin Zone, except along 2 ($d$-wave), 4 ($g$-wave) or 6 ($i$-wave) nodal surfaces Šmejkal et al. (2020); Mazin et al. (2021), leading to various spintronic effects Šmejkal et al. (2020); Mazin et al. (2021); Feng et al. (2022); Reichlova et al. (2024); Šmejkal et al. (2022c); Bai et al. (2024); Naka et al. (2019); Shao et al. (2021).
The theoretically predicted and experimentally confirmed altermagnets are hitherto characterized by an even number of magnetic sublattices, whose alternating magnetic moments are accompanied by an alternating orientation of the atomic site environment in direct space Šmejkal et al. (2022a, b); Guo et al. (2023); Bai et al. (2024). In a typical $d$-wave altermagnet, the alternating magnetic sublattices are related by a $180^{\circ}$ spin rotation combined with a lattice rotation (proper or improper). This alternation of magnetic sites is commonly achieved by the anisotropic crystal environment of the magnetic sublattices Šmejkal et al. (2022a, 2020), but can also be achieved by electronic correlations Leeb et al. (2024); Dürrnagel et al. (2024).
In this letter we address the ferroic nature of altermagnets, by studying the partial-wave expansion of their direct-space spin density. Previous reports have shown that some altermagnetic candidates host ferroically ordered magnetic multipoles on the magnetic sublattices Bhowal and Spaldin (2024); Fernandes et al. (2024); Verbeek et al. (2024). Both approaches aim to describe the anisotropy of the spin density beyond the dipolar approximation ($s$-wave), and, while being complementary, they are also fundamentally distinct. The magnetic multipoles approach is inspired in classical electromagnetism, while the partial-waves analysis arises from atomic physics. We show that the use of partial waves reveals a direct correlation between the anisotropy of the atomic spin density in direct-space, and the spin-splitting of the electronic band structure in momentum space.
Here we demonstrate the presence of a ferroically ordered $g$-wave form factor of the atomic spin density in CrSb and MnTe, which we consider here to be canonical altermagnets. We show that the ferroically ordered $g$-wave spin density coexists with an antiferroic (antiparallel) order of the atomic magnetic dipole moments (Néel order). In A2O3M2Se2 (A= La, Sr and M= Mn, Fe, Co) with a Lieb lattice structure formed by the M and O sites, we demonstrate a mixture of, in the one hand, a ferroically ordered $d$-wave spin density with antiferroic magnetic dipoles at the M sites, and in the other hand a pure $d$-wave atomic spin density (no magnetic dipole) around the O site.
As a key result, we further propose a pure form of atomic altermagnetism, i.e., a direct-space spin density without any dipolar moments at any atomic site but with an on-site ferroically ordered $d$-, $g$- or $i$-wave form factor, thus enabling the realization of altermagnetism without the need to have the staggered Néel order of local dipole moments. In addition, we show that strong correlations can conspire with orbital states to generate an on-site altermagnetic $d$-wave spin density, without being promoted by the symmetry of the ionic-crystal lattice. We demonstrate such an atomic altermagnetic state both in the non-relativistic limit and in the presence of the relativistic spin-orbit coupling (SOC) in the Ba2CaOsO6 material Maharaj et al. (2020); Paramekanti et al. (2020); Voleti et al. (2021).
It is well established that $d$-wave altermagnets can host spin-dependent conductivities, leading to giant-magnetoresistance (GMR) Šmejkal et al. (2022c) and spin-splitter effects Naka et al. (2019); González-Hernández et al. (2021); Bose et al. (2022). In the latter case, a pure spin current flows transverse to the applied electric bias. We calculate a large spin-splitter response in both A2O3M2Se2  and Ba2CaOsO6, demonstrating that the altermagnetic spin density on a single atomic site without Néel order, i.e., a pure atomic altermagnet (in the case of Ba2CaOsO6), can generate spin-splitter effect.
Figure 2: DFT calculation for La2O3Mn2Se2, without (a-c) and with (d-j) SOC. (a) and (d) Band structure with $S_{z}$ spin polarization showed in color-scale. (b) Iso-surface of the non-relativistic spin density. (c) Energy iso-surface cut in the plane $k_{z}=0$. (e-g) Iso-surface of the spin density (with SOC) from different views, projected along the spin components $[001]$ (e), $[110]$ (f) and $[\bar{1}10]$(g). Next to each spin component we highlight the ferroically ordered partial-wave of the spin density: $d_{x^{2}-y^{2}}$ (e), $d_{zx}-d_{yz}$(f) and $d_{zx}+d_{yz}$ (g). Iso-energy lines on planes that contain the $\Gamma$ point and are normal to the directions $[001]$ (h), $[110]$ (i) and $[\bar{1}10]$(j), with the color scale representing the spin polarization on the axis normal to the planes.
Partial-wave expansion of the spin density – The spin density, as a property of the Fermi sea, can be defined as $\mathbf{S}(\mathbf{r})=\frac{\hbar}{2}\sum_{\mathbf{k},n}f_{n\mathbf{k}}\psi_{n\mathbf{k}}^{\dagger}(\mathbf{r})\bm{\sigma}\psi_{n\mathbf{k}}(\mathbf{r})$, where $\psi_{n\mathbf{k}}(\mathbf{r})$ is the spinorial eigenfunction of wave-vector $\mathbf{k}$ and band index $n$, $f_{n\mathbf{k}}$ is the occupation number and $\bm{\sigma}$ are the Pauli matrices representing the spin degree of freedom. We define $\mathbf{S}_{\alpha}(\mathbf{r})=\mathbf{S}(\mathbf{r}-\mathbf{r}_{\alpha})$ as the spin density with respect to an atomic site $\mathbf{r}_{\alpha}$, and we further expand each spin component in spherical harmonics as ${S}^{i}_{\alpha}(\mathbf{r})=\sum_{lm}S^{i}_{\alpha,lm}(r)Y_{lm}(\theta,\phi)$ where $Y_{lm}(\theta,\phi)$ are real spherical harmonics Blanco et al. (1997) and $(r,\theta,\phi)$ denote the spherical coordinates of $\mathbf{r}-\mathbf{r}_{\alpha}$. The magnetic dipole approximation allows us to define an on-site magnetic dipole as the radial integral of the first term in the expansion $S^{i}_{\alpha,00}(r)$. The magnetic dipoles alone, however, are unable to capture the ferroic nature of altermagnetism Bhowal and Spaldin (2024); Fernandes et al. (2024).
Canonical altermagnets – In a canonical altermagnet, such as CrSb (Type Ia in Fig. 1a,d), the magnetic dipoles align antiferroically, but the spin density around each magnetic site is anisotropic. For CrSb there is a ferroically ordered contribution with the symmetries of a $g_{\bar{3}}\equiv Y_{4,-{3}}$ harmonic (See Fig. 1d). Remarkably, unlike conventional multipolar orders which originate from SOC Santini and Amoretti (2000); Zhao et al. (2016), the ferroic order in the above altermagnets is stabilized by the antiferroic magnetic dipoles distributed in the crystal according to the altermagnetic spin group symmetry Šmejkal et al. (2022a). The $g_{\bar{3}}$ harmonic has the same nodal structure as the spin splitting in the electronic band structure, which results from the spin group of CrSb containing operations such as $[C_{2}||C_{6z}|\vec{\tau}]$ combining a two-fold spin rotation with a six-fold lattice rotation, and a half unit cell translation along the z-axis as marked in Fig. 1a. In the Supplementary Information (SI), we show similar analysis for MnTe.
Lieb lattice candidate – Next we discuss the A2O3M2Se2 family (with special focus on A=La, M=Mn). The crystal structure is body centered tetragonal with space group $I4/mmm$ (No. 139) Wei et al. (2024); Landsgesell et al. (2013); Free et al. (2011). The Mn atoms are located at the Wyckoff position $4c$, and together with the oxygen ligands located at Wyckoff position $2b$, they form a layered structure that resembles the Lieb lattice Lieb (1989); Inui et al. (1994).
The total spin density obtained by density-functional theory (DFT), without spin-orbit coupling, is shown in Fig. 1b. While La2O3Mn2Se2 is a Mott insulator, the symmetries of the spin density in both direct and momentum space can be captured by DFT. We decompose this spin density into spherical harmonics around Mn and O sites. In the Mn site, besides the $s$-wave staggered magnetic dipoles (see top panel in Fig 1e) we identify two ferroically ordered contributions with the symmetries of the $d_{x^{2}-y^{2}}$ and $g_{2}\equiv Y_{4,2}$ harmonics. Both partial waves have the same $d$-wave nodal structure, and they deform the otherwise spherically symmetric spin density. In the Oxygen site, which has no magnetic dipole moment, we see that its atomic spin density is fully described by a $d_{x^{2}-y^{2}}$ form factor. In the bottom panel of Fig. 1e we show the $d$-wave direct-space order parameter that is obtained by adding up the $d_{x^{2}-y^{2}}$ form factors on MnA, MnB and O sublattices. Indeed, the spin density in the material follows the spin point group ${}^{2}4/^{1}m^{1}m^{2}m$, which classifies it as a $d$-wave altermagnet. The electronic band structure is spin-split (see Fig. 2a), with spin splitting reaching a 1 eV scale, and the spin polarization of the energy iso-surfaces in momentum space (Fig. 2c) shows the same $d$-wave character as the ferroically ordered $d_{x^{2}-y^{2}}$ projection of the spin density in direct space (Fig. 2b).
A key factor behind the large splitting in the A2O3M2Se2 family is the Lieb lattice structure. Here, the anisotropy of the exchange enters at the second Mn-Mn neighbor. If we consider the MnA sublattice, the links with neighbors at $(a,0,0)$ and $(0,a,0)$ are different, because the first one contains the O-site in the middle. This anisotropic exchange at close distance can lead to large spin splitting and strength of the altermagnetic order.
We emphasize that, in the non-relativistic collinear regime, both the direct-space spin density and the momentum-space spin polarization of the band structure have spin degenerate nodal planes orthogonal to the directions $[110]$ and $[\bar{1}10]$. When SOC is included in the DFT calculations (Néel order along $[001]$), the nodal planes show a weak $d$-wave spin-splitting, with spin polarization normal to the respective plane. Thus, when looking at the plane in momentum space normal to the $[110]$ axis that crosses the $\Gamma$ point (Fig. 2i), the component $S_{110}$ shows a $d$-wave character, with nodal planes normal to the $[\bar{1}10]$ and $[001]$ axes. The corresponding $S_{110}$ component of the direct-space spin density displays the same $d$-wave nodal structure (Fig. 2f). Analogously, if we focus on the plane normal to the $[\bar{1}10]$ axis, both the direct-space spin density (Fig. 2g), and the spin polarization of the energy iso-surface (Fig. 2j) have nodal planes normal to the $[110]$ and $[001]$ axes. Thus, there is a clear correspondence between the nodal structure of the direct-space spin density, and the spin polarization of the band structure in momentum space. More broadly, as we show below, each spin component displays a different d-wave form-factor, consistent with the analysis of Ref. Fernandes et al. (2024). A full characterization of the partial-wave decomposition of the direct space spin density for La2O3Mn2Se2 can be seen in the SI.
In La2O3Mn2Se2, the $[110]$ ($[\bar{1}10]$) component of the spin density is two orders of magnitude smaller than the $S_{001}$ component. Additional insight into the spin ordering in the presence of SOC is provided by the following minimal model:
$\begin{split}h(\mathbf{k})=&\lambda_{1}(k_{x}^{2}-k_{y}^{2})\sigma_{z}+\\ &\lambda_{2}(k_{y}-k_{x})k_{z}\sigma_{[110]}+\lambda_{2}(k_{x}+k_{y})k_{z}\sigma_{[\bar{1}10]}\;,\end{split}$ (1)
where $\mathbf{k}$ is the wave-vector and $\sigma_{z}$, $\sigma_{[110]}$, and $\sigma_{[\bar{1}10]}$ are defined as $\bm{\sigma}\cdot\mathbf{v}$, with $\mathbf{v}$ a unit vector along $[001]$, $[110]$ and $[\bar{1}10]$ respectively. Without SOC, $\lambda_{2}=0$, and when SOC is included the additional $d$-wave splitting is captured by $\lambda_{2}$, with $\lambda_{2}<<\lambda_{1}$. An important consequence of the existence of two nodal planes for each of the three independent spin directions is that the band structure will only be spin-degenerate along nodal lines in the presence of SOC rather than nodal planes.
Pure atomic altermagnet candidate – We consider now Ba2CaOsO6, whose crystal structure is face-centered cubic with space group $Fm\bar{3}m$ (No. 225) Thompson et al. (2014). The Os atom occupies the Wyckoff position $4a$. This material has drawn significant attention owing to unexpected oscillations in the zero-field muon spin relaxation Thompson et al. (2014), which suggest a phase transition with TRSB at $T\sim 50\;K$. Based on neutron spectroscopy and $X$-ray diffraction measurements, it was proposed that this phase transition can be explained by a ferro-octupolar phase Maharaj et al. (2020), with magnetic octupoles ordered on the Os $5d^{2}$ electrons. Later on, this claim was supported from an ab initio perspective Fiore Mosca et al. (2022), using VASP and on-site density matrix control Allen and Watson (2014). To our knowledge, however, little has been said about the connection between this ferro-octupolar phase and the spin splitting in the electronic band structure.
Despite SOC being strong in this double perovskite, it is still instructive to explore its non-relativistic limit, because also in this limit we find a correspondence between the direct and momentum space order parameters. We calculate the spin density from DFT, including correlations on the Os site using the Liechtenstein’s spherically symmetric DFT+U approach Liechtenstein et al. (1995), with parameters $U=3.2$ eV and $J=0.5$ eV, in agreement with the pioneer study on the compound Fiore Mosca et al. (2022). Without SOC, the collinear spin density around the Os site is shown in Fig. 1c. Below, in Fig. 1f we see that the harmonic decomposition is given by $d_{xy}$ and $g_{\bar{2}}\equiv Y_{4,-2}$. Both spherical harmonics have $d$-wave symmetry in the $xy$ plane, but unlike La2O3Mn2Se2, the nodal planes are normal to $[100]$ and $[010]$. There is no magnetic dipole moment, as can be seen in the top panel of Fig. 1f, making this material a pure atomic altermagnet. The corresponding iso-energy surfaces (Fig. 3c) follow the $[C_{2}||C_{4z}]$ spin symmetry.
In Fig. 3a we show the band structure along a path in the $(k_{x},k_{y})$-plane that bisects the two nodal planes. We identify a spin splitting of 0.2 eV just below the Fermi level. We emphasize that, because Ba2CaOsO6 is a Mott insulator, the bands crossing the Fermi level in Fig. 3a will be pushed away by the on-site repulsion. Further investigation of the impact of strong correlations on the band structure is, however, beyond the scope of our work, which focuses on the symmetry-enforced shape of the spin density on the Os site. In this regard, we note that the $d_{xy}$ form factor of the spin density obtained via DFT+U is consistent with the strong-coupling perspective on the octupolar moment of Os $5d^{2}$, which can be understood as the $z$-component of the pseudo-spin that characterizes the non-Krammers doublet ground-state of the Os atom, arising from the combined effects of crystal fields and SOC on the localized $J=2$ state Paramekanti et al. (2020); Voleti et al. (2021).
Figure 3: DFT+U calculation for Ba2CaOsO6 without (a-c) and with (d-j) SOC . (a), (d) Band structure colored according to the $S_{z}$ spin polarization. (b) Iso-surface of the direct-space spin density (no SOC) around the Os site. (c) Energy iso-surface cut in the plane $k_{z}=0$. Iso-surface of the spin density around the Osmium site (with SOC), projected along the spin components $[100]$ (e), $[010]$ (f) and $[001]$(g). Energy iso-surface cuts in planes that contain the $\Gamma$ point and are normal to directions $[100]$ (h), $[010]$ (i) and $[001]$(j), with color scale representing the spin polarization along the axis normal to the planes.
Once SOC is included, the spin density in Ba2CaOsO6, in agreement with previous reports Fiore Mosca et al. (2022), becomes strongly noncoplanar, with equivalent $d$-wave form factors along the $x$, $y$, and $z$ axes. These results are consistent with the analysis of Ref. Fernandes et al. (2024), which found that the inclusion of SOC makes the spin-density noncoplanar in some altermagnets. Figure 3d shows the band structure (with SOC) along a path in the $x-y$ plane, bisecting the planes normal to $[100]$ and $[010]$ axes. We also show the $z$-component of the spin polarization in a color scale. Note that, in the presence of SOC, the system displays a band gap of around 1 eV. We note, however, that strong correlations will render the material a Mott insulator.
The $x$, $y$ and $z$ components of the spin density around the Os atom are shown in Figs. 3 e-g. Below each spin density, in Figs. 3 h-j, we show energy iso-surface cuts with the spin polarization (component normal to each plane) plotted in color scale. The spin density can be understood as 3 orthogonal $d$-wave form factors corresponding to the $x$, $y$ and $z$ spin components. The nodal structure of the spin density is reflected in the spin polarization of the energy iso-surfaces. In analogy with Eq. 1, a simple model capturing these noncoplanar altermagnetic $d$-wave form factors is given by Fernandes et al. (2024):
$h(\mathbf{k})=\lambda k_{x}k_{y}\sigma_{z}+\lambda k_{z}k_{x}\sigma_{y}+\lambda k_{y}k_{z}\sigma_{x}\text{ .}$ (2)
Note that, because of the cubic symmetry of the crystal structure, and in contrast to La2O3Mn2Se2, all spin components have the same prefactor $\lambda$.
The emergence of the atomic $d$-wave form factors in Ba2CaOsO6, with or without SOC, comes from an interplay between correlations and orbital ordering. The on-site occupation matrix of the $d$ orbitals captures the $d$-wave character Fiore Mosca et al. (2022), which then gets imprinted in both the direct-space spin density and the spin splitting in the band structure.
Figure 4: Estimation of the spin-splitter angle for the non-relativistic calculation of both La2O3Mn2Se2 (a) and Ba2CaOsO6 (b). See in Supplementary Information the calculation of the GMR coefficients.
Spin currents – We calculate the magnitude of the spin currents in the non-relativistic limit for both La2O3Mn2Se2 and Ba2CaOsO6, by quantifying the spin-splitter angle $\alpha$, which measures the efficiency in the charge-to-spin conversion, and the GMR coefficient Šmejkal et al. (2022c). Our results are shown in Fig. 4, where the spin-splitter angle is plotted as a function of the Fermi energy. In the explored energy interval, the spin-splitter angle for Ba2CaOsO6 (Fig. 4a) reaches a maximum of 26 , while for La2O3Mn2Se2 it is even higher, reaching 44. The GMR coefficient (see SI) in La2O3Mn2Se2 reaches 900% for energies close to the Fermi energy, which is one order of magnitude higher than the 100% scale of the initial altermagnetic GMR prediction, considering RuO2 Šmejkal et al. (2022c). While La2O3Mn2Se2 is not metallic, we expect the spin-splitter angle and the GMR coefficient to be equally large in related metallic altermagnetic candidates on layered Lieb lattices, such as UCr2Si2Lemoine et al. (2018), KV2Se2O Jiang et al. or RbV2Se2O Zhang et al. (2024). In addition to the $d$-wave form factor in the sites hosting the magnetic dipoles, the $d$-wave atomic spin density will be present, without dipole, at the O site in RbV2Se2O and KV2Se2O, and at the U site at UCr2Si2C. In the SI, we show calculations of the electronic structure for UCr2Si2C, and we comment as well on the nature of the spin conductivity tensor in Ba2CaOsO6(with SOC), when the spin density is described by three noncoplanar $d$-wave form factors.
Discussion – While the notion of multipolar magnetism is naturally connected with altermagnetism Bhowal and Spaldin (2024); Fernandes et al. (2024), one should be careful when comparing these two concepts. Not every direct-space ferroically ordered magnetic multipole, corresponds to the collinear $d$-, $g$-, or $i$-wave altermagnetic order with the characteristic anisotropic nodal spin polarization in the momentum-space electronic structure. As an illustrative example, consider Mn3Sn, which has a phase between $T_{N2}=50$K and $T_{N1}=420$ K in which Mn magnetic moments form an inverse triangular spin structure. Although the material exhibits octupolar order $T_{x}^{\gamma}$ Suzuki et al. (2017), its magnetic order on the crystal is non-collinear and, therefore, is not altermagnetic. Correspondingly, the resulting spin polarization in the momentum space takes noncollinear and nodeless form Hellenes et al. (2023).
Conclusions— In this work we extend the known mechanisms that induce altermagnetism, to include pure atomic altermagnetism. The interplay between strong correlations and orbital occupation (possibly with strong SOC) generates a spin density which lacks magnetic dipole moments, while having a $d$-wave component which is also reflected in the corresponding spin-polarization symmetry of the electronic structure in the momentum space. In La2O3Mn2Se2 we demonstrate the coexistence of dipole and higher-order partial-wave components on the Mn-sites, with the absence of the dipole and the presence of the higher-order partial-wave component (atomic altermagnetism) on the O-sites. We estimate the spin-splitter angle and the GMR coefficient in this material as 44 degrees and 900%, respectively.
We identified pure atomic altermagnetism with $d$-wave symmetry in the non-relativistic limit of Ba2CaOsO6. When including SOC, the spin density has the structure of three noncoplanar (and orthogonal) $d$-waves form factors. In all cases, we identify one-to-one correspondence between the ferroically ordered $d$-wave harmonic of the spin density in direct space, and the angular dependence of the spin splitting around the $\Gamma$ point.
[…]
Acknowledgments— RJU, VKB, RZ, JS and LS acknowledge funding by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation)- TRR 173– 268565370 (project A03) and TRR 288– 422213477 (project A09 and B05). TJ acknowledges Ministry of Education of the Czech Republic Grant CZ.02.01.01/00/22008/0004594 and ERC Advanced Grant no. 101095925. R.M.F. was supported by the Air Force Office of Scientific Research under Award No. FA9559-21-1-0423.

### v2

@@
 Abstract
-Altermagnetism has been recently experimentally verified by photoemission mapping of the spin order in momentum space in MnTe and CrSb, which feature two anisotropic sublattices with antiparallel magnetic dipole moments.
+Altermagnetism has been recently verified experimentally by photoemission mapping of the spin order in momentum space in MnTe and CrSb, which feature two anisotropic sublattices with antiparallel magnetic dipole moments.
 In this work, we explicitly demonstrate the presence of an even-parity ferroically ordered non-dipolar spin density on atomic sites, i.e.
-atomic altermagnetism, in MnTe, La2O3Mn2Se2 and Ba2CaOsO6.
+atomic altermagnetism, in MnTe, KV2Se2O and Ba2CaOsO6.
 We do so through spin-symmetry analysis and partial-wave decomposition of the spin density obtained by first-principles calculations.
 In MnTe we show a ferroically ordered $g$-wave form factor in the spin density around the Mn site.
-In the A2O3M2Se2 family (A= La, Sr and M= Mn, Fe, Co), we show that there is a ferroically ordered $d$-wave form factor coexisting with the antiferroic magnetic dipoles in the M site, while the O site shows no dipole but a pure $d$-wave atomic spin density.
+In KV2Se2O (and related Lieb lattice compounds), we show that there is a ferroically ordered $d$-wave form factor coexisting with the antiferroic magnetic dipoles in the V site, while the O site shows no dipole but a pure $d$-wave atomic spin density.
 In the Mott-insulating Ba2CaOsO6, as a key result, we reveal a pure form of atomic altermagnetism – absent of any dipolar sublattice order.
 This highlights that the altermagnetic order can exist without a Néel vector formed by antiferroic dipole moments on an even number of crystal sublattices, underlining its distinction from collinear Néel antiferromagnetic order.
-Our calculations predict that La2O3Mn2Se2 and Ba2CaOsO6 can exhibit giant spin-splitter angles of up to 42° and 26° respectively, thus demonstrating the possibility of large altermagnetic responses without requiring the staggered Néel order of local dipole moments.
-Figure 1: Partial-wave expansion of the spin density from non-relativistic DFT calculations on CrSb, A2O3M2Se2, and Ba2CaOsO6.
-(d), (e) and (f) show the radial dependence of the relevant partial-wave components $S_{lm}(r)$ as defined in the main text.
-(a) CrSb unit cell, where black arrows represent the magnetic dipoles and, superimposed, we show the ferroically ordered $g_{\bar{3}}=Y_{4,-3}$ harmonic.
-(d) Radial dependence of the order parameters.
-(top) Antiferroic $s$-wave on each sublattice, (bottom) ferroic $g$-wave form factor.(b) Unit cell of La2O3Mn2Se2 with an iso-surface of the spin density.
-Mn sites have opposite magnetic dipole moments, and O site has no dipole but a $d_{x^{2}-y^{2}}$ spin density.
-(e) (top) $s$-wave harmonic expansion centered at each Mn site.
-(bottom) $d$-wave direct-space order parameter, formed by adding up the $d_{x^{2}-y^{2}}$ contributions of MnA, MnB and O sublattices.
-(c) Total spin density iso-surface around the Os site in Ba2CaOsO6 within its octahedral environment.
-(f) shows the respective relevant partial-waves, with zero $s$-wave contribution (no magnetic dipole) and two ferroic $d$-wave harmonics $d_{xy}$ and $g_{\bar{2}}\equiv Y_{4,-2}$.
+Our calculations predict that KV2Se2O and Ba2CaOsO6 can exhibit giant spin-splitter angles of up to 42° and 26° respectively, thus demonstrating the possibility of large altermagnetic responses without requiring the staggered Néel order of local dipole moments.
+Figure 1: Partial-wave expansion of the spin density from non-relativistic DFT for CrSb, V2Te2O, and Ba2CaOsO6.
+Panels (d–f) show the radial profiles of the key partial-wave coefficients $S_{lm}(r)$: top panels display antiferroic $s$-wave (dipolar) components; bottom panels show ferroically ordered higher harmonics.
+(a) CrSb unit cell with antiferroic dipoles (black arrows) and ferroic $g_{\bar{3}}\equiv Y_{4,-3}$.
+(b) V2Te2O unit cell and spin density iso-surface; V sites show antiferroic dipoles and ferroic $d_{x^{2}-y^{2}}$ and $g_{2}\equiv Y_{42}$ components; the O site hosts a pure $d_{x^{2}-y^{2}}$ density.
+(c) Ba2CaOsO6 spin density iso-surface, showing no dipole but a pure $d$-wave spin density
 Altermagnetism is a recently identified new type of magnetism characterized by spin order in both direct and momentum space with $d$-, $g$-, or $i$-wave symmetry Šmejkal et al.
@@
 (2022a).
-These spin symmetries lead to an unconventional complex altermagnetic spin density Šmejkal et al.
+These symmetries lead to unconventional spin densities Šmejkal et al.
 (2020) and exchange fields Šmejkal et al.
@@
 (2024).
-Recently, altermagnetism has been experimentally confirmed by momentum-space and direct-space photoemission spectroscopy and microscopy in representative binary-compounds MnTe and CrSb with ordering above room temperature Krempaskỳ et al.
+Altermagnetism has been experimentally confirmed via photoemission spectroscopy and microscopy, in the binary-compounds MnTe and CrSb with ordering above room temperature Krempaský et al.
 (2024); Lee et al.
@@
 (2024).
-The discovery of the altermagnetic spin symmetry class also provided a unifying explanation of the previous reports of unconventional time-reversal symmetry breaking (TRSB) in the electronic structure and related effects, such as the anomalous Hall effect Šmejkal et al.
+More recently, two room-temperature metallic $d$-wave altermagnets, KV2Se2Jiang et al.
+(2025) and Rb1-δV2Te2O Zhang et al.
+(2025), have been probed with neutron magnetic resonance, and spin resolved photo-emission measurements, respectively.
+The discovery of the altermagnetic spin symmetry class also provided a unifying explanation of the previous reports of unconventional time-reversal symmetry breaking (TRSB), such as the anomalous Hall effect Šmejkal et al.
 (2020); Mazin et al.
@@
 (2024).
-In a typical $d$-wave altermagnet, the alternating magnetic sublattices are related by a $180^{\circ}$ spin rotation combined with a lattice rotation (proper or improper).
 This alternation of magnetic sites is commonly achieved by the anisotropic crystal environment of the magnetic sublattices Šmejkal et al.
-(2022a, 2020), but can also be achieved by electronic correlations Leeb et al.
+(2022a, 2020); Jaeschke-Ubiergo et al.
+(2024); Gomonay et al.
+(2024), but can also be achieved by electronic correlations Leeb et al.
 (2024); Dürrnagel et al.
+(2024); Giuli et al.
 (2025).
-In this letter we address the ferroic nature of altermagnets, by studying the partial-wave expansion of their direct-space spin density.
-Previous reports have shown that some altermagnetic candidates host ferroically ordered magnetic multipoles on the magnetic sublattices Bhowal and Spaldin (2024); Fernandes et al.
-(2024); Verbeek et al.
+Previous reports have suggested that altermagnetism is linked to ferroically ordered magnetic multipoles on the magnetic sublattices Bhowal and Spaldin (2024); Fernandes et al.
+(2024); McClarty and Rau (2023); Verbeek et al.
+(2024); Schiff et al.
 (2025).
-Both approaches aim to describe the anisotropy of the spin density beyond the dipolar approximation ($s$-wave), and, while being complementary, they are also fundamentally distinct.
-The magnetic multipoles approach is inspired in classical electromagnetism, while the partial-waves analysis arises from atomic physics.
-We show that the use of partial waves reveals a direct correlation between the anisotropy of the atomic spin density in direct-space, and the spin-splitting of the electronic band structure in momentum space.
-Here we demonstrate the presence of a ferroically ordered $g$-wave form factor of the atomic spin density in CrSb and MnTe, which we consider here to be canonical altermagnets.
-We show that the ferroically ordered $g$-wave spin density coexists with an antiferroic (antiparallel) order of the atomic magnetic dipole moments (Néel order).
-In A2O3M2Se2 (A= La, Sr and M= Mn, Fe, Co) with a Lieb lattice structure formed by the M and O sites, we demonstrate a mixture of, in the one hand, a ferroically ordered $d$-wave spin density with antiferroic magnetic dipoles at the M sites, and in the other hand a pure $d$-wave atomic spin density (no magnetic dipole) around the O site.
-As a key result, we further propose a pure form of atomic altermagnetism, i.e., a direct-space spin density without any dipolar moments at any atomic site but with an on-site ferroically ordered $d$-, $g$- or $i$-wave form factor, thus enabling the realization of altermagnetism without the need to have the staggered Néel order of local dipole moments.
-In addition, we show that strong correlations can conspire with orbital states to generate an on-site altermagnetic $d$-wave spin density, without being promoted by the symmetry of the ionic-crystal lattice.
-We demonstrate such an atomic altermagnetic state both in the non-relativistic limit and in the presence of the relativistic spin-orbit coupling (SOC) in the Ba2CaOsO6 material Maharaj et al.
+In this work we aim to quantify the altermagnetic direct-space order parameter Fernandes et al.
+(2024); McClarty and Rau (2023), by studying the partial-wave expansion of the spin density obtained from Density Funtional Theory (DFT).
+We demonstrate the presence of a ferroically ordered $g$-wave form factor of the atomic spin density in CrSb and MnTe, which we consider here to be canonical altermagnets.
+In a series of compounds with a Lieb lattice structure Mazin et al.
+(2023); Kaushal and Franz (2024) (e.g., KV2Se2Jiang et al.
+(2025)), we find a ferroic $d$-wave spin density which is comparable in size with the antiferroic dipoles on V sites, and a pure $d$-wave spin density (no dipole) on the O site.
+As a key result, in Ba2CaOsO6 Maharaj et al.
 (2020); Paramekanti et al.
 (2020); Voleti et al.
-(2021).
-It is well established that $d$-wave altermagnets can host spin-dependent conductivities, leading to giant-magnetoresistance (GMR) Šmejkal et al.
-(2022c) and spin-splitter effects Naka et al.
-(2019); González-Hernández et al.
-(2021); Bose et al.
-(2022).
-In the latter case, a pure spin current flows transverse to the applied electric bias.
-We calculate a large spin-splitter response in both A2O3M2Se2  and Ba2CaOsO6, demonstrating that the altermagnetic spin density on a single atomic site without Néel order, i.e., a pure atomic altermagnet (in the case of Ba2CaOsO6), can generate spin-splitter effect.
-Figure 2: DFT calculation for La2O3Mn2Se2, without (a-c) and with (d-j) SOC.
-(a) and (d) Band structure with $S_{z}$ spin polarization showed in color-scale.
+(2021), we propose a pure form of atomic altermagnetism, i.e., a direct-space spin density without any dipolar moments at any atomic site but with an on-site ferroically ordered $d$-wave form factor, thus enabling the realization of altermagnetism without the need to have the staggered Néel order of local dipole moments.
+We show that correlations and orbital ordering can generate $d$-wave altermagnetic spin density, without being promoted by particular crystal symmetry.
+We also find large spin-splitter responses in the nonrelativistic limits of KV2Se2O and Ba2CaOsO6, demonstrating that pure atomic altermagnetism— in the case of Ba2CaOsO6, without Néel order—can lead to large altermagnetic responses.
+Figure 2: DFT results for KV2Se2O without (a–c) and with (d–j) SOC.
+(a,d) Band structure with $S_{z}$ spin polarization shown in color scale.
 (b) Iso-surface of the non-relativistic spin density.
-(c) Energy iso-surface cut in the plane $k_{z}=0$.
-(e-g) Iso-surface of the spin density (with SOC) from different views, projected along the spin components $[001]$ (e), $[110]$ (f) and $[\bar{1}10]$(g).
-Next to each spin component we highlight the ferroically ordered partial-wave of the spin density: $d_{x^{2}-y^{2}}$ (e), $d_{zx}-d_{yz}$(f) and $d_{zx}+d_{yz}$ (g).
-Iso-energy lines on planes that contain the $\Gamma$ point and are normal to the directions $[001]$ (h), $[110]$ (i) and $[\bar{1}10]$(j), with the color scale representing the spin polarization on the axis normal to the planes.
-Partial-wave expansion of the spin density – The spin density, as a property of the Fermi sea, can be defined as $\mathbf{S}(\mathbf{r})=\frac{\hbar}{2}\sum_{\mathbf{k},n}f_{n\mathbf{k}}\psi_{n\mathbf{k}}^{\dagger}(\mathbf{r})\bm{\sigma}\psi_{n\mathbf{k}}(\mathbf{r})$, where $\psi_{n\mathbf{k}}(\mathbf{r})$ is the spinorial eigenfunction of wave-vector $\mathbf{k}$ and band index $n$, $f_{n\mathbf{k}}$ is the occupation number and $\bm{\sigma}$ are the Pauli matrices representing the spin degree of freedom.
-We define $\mathbf{S}_{\alpha}(\mathbf{r})=\mathbf{S}(\mathbf{r}-\mathbf{r}_{\alpha})$ as the spin density with respect to an atomic site $\mathbf{r}_{\alpha}$, and we further expand each spin component in spherical harmonics as ${S}^{i}_{\alpha}(\mathbf{r})=\sum_{lm}S^{i}_{\alpha,lm}(r)Y_{lm}(\theta,\phi)$ where $Y_{lm}(\theta,\phi)$ are real spherical harmonics Blanco et al.
+(c) Energy iso-surface in the $k_{z}=0$ plane.
+(e–g) Spin density iso-surfaces with SOC, projected along $[001]$, $[110]$, and $\bar{1}10]$, respectively, showing ferroic partial-wave components: $d_{x^{2}-y^{2}}$, $d_{zx}-d_{yz}$, and $d_{zx}+d_{yz}$, respectively.
+(h–j) Iso-energy lines in planes normal to $[001]$, $[110]$, and $\bar{1}10]$, colored by spin polarization normal to each plane.
+Partial-wave expansion of the spin density—The spin density can be defined as $\mathbf{S}(\mathbf{r})=\frac{\hbar}{2}\sum_{\mathbf{k},n}f_{n\mathbf{k}}\psi_{n\mathbf{k}}^{\dagger}(\mathbf{r})\bm{\sigma}\psi_{n\mathbf{k}}(\mathbf{r})$, where $\psi_{n\mathbf{k}}(\mathbf{r})$ is the spinorial eigenfunction of wave-vector $\mathbf{k}$ and band index $n$, $f_{n\mathbf{k}}$ is the occupation number and $\bm{\sigma}$ are the Pauli matrices representing the spin degree of freedom.
+We define $\mathbf{S}_{\alpha}(\mathbf{r})=\mathbf{S}(\mathbf{r}-\mathbf{r}_{\alpha})$ as the spin density with respect to an atomic site $\mathbf{r}_{\alpha}$, and we further expand each spin component in partial waves as ${S}^{i}_{\alpha}(\mathbf{r})=\sum_{lm}S^{i}_{\alpha,lm}(r)Y_{lm}(\theta,\phi)$ where $Y_{lm}(\theta,\phi)$ are real spherical harmonics Blanco et al.
 (1997) and $(r,\theta,\phi)$ denote the spherical coordinates of $\mathbf{r}-\mathbf{r}_{\alpha}$.
@@
 (2024).
-Canonical altermagnets – In a canonical altermagnet, such as CrSb (Type Ia in Fig.
+Canonical altermagnets—In a canonical altermagnet, such as CrSb (Type Ia, see Fig.
 1a,d), the magnetic dipoles align antiferroically, but the spin density around each magnetic site is anisotropic.
@@
 1d).
-Remarkably, unlike conventional multipolar orders which originate from SOC Santini and Amoretti (2000); Zhao et al.
+Remarkably, unlike conventional multipolar orders which originate from spin-orbit coupling (SOC) Santini and Amoretti (2000); Zhao et al.
 (2016), the ferroic order in the above altermagnets is stabilized by the antiferroic magnetic dipoles distributed in the crystal according to the altermagnetic spin group symmetry Šmejkal et al.
 (2022a).
-The $g_{\bar{3}}$ harmonic has the same nodal structure as the spin splitting in the electronic band structure, which results from the spin group of CrSb containing operations such as $[C_{2}||C_{6z}|\vec{\tau}]$ combining a two-fold spin rotation with a six-fold lattice rotation, and a half unit cell translation along the z-axis as marked in Fig.
+The $g_{\bar{3}}$ harmonic has the same nodal structure as the spin splitting in the electronic band structure, which results from the spin group of CrSb containing operations such as $C_{2}||C_{6z}|\vec{\tau}]$ combining a two-fold spin rotation with a six-fold lattice rotation, and a half unit cell translation along the z-axis as marked in Fig.
 1a.
 In the Supplementary Information (SI), we show similar analysis for MnTe.
-Lieb lattice candidate – Next we discuss the A2O3M2Se2 family (with special focus on A=La, M=Mn).
-The crystal structure is body centered tetragonal with space group $I4/mmm$ (No.
-139) Wei et al.
+Lieb lattice candidates—Next we discuss a series of materials whose crystal structure can be understood as stacked layers of Lieb lattices Mazin et al.
+(2023).
+This list includes the quasi 2D monolayers of V2A2O (A=Se, Te) Lin et al.
+(2018); Qi et al.
+(2024), and the bulk structures KV2Se2Jiang et al.
+(2025), A2O3M2Se2 (A=La, M=Mn, Fe) Wei et al.
 (2025); Landsgesell et al.
 (2013); Free et al.
+(2011) and UCr2Si2Lemoine et al.
 (2018).
-The Mn atoms are located at the Wyckoff position $4c$, and together with the oxygen ligands located at Wyckoff position $2b$, they form a layered structure that resembles the Lieb lattice Lieb (1989); Inui et al.
-(1994).
-The total spin density obtained by density-functional theory (DFT), without spin-orbit coupling, is shown in Fig.
-1b.
-While La2O3Mn2Se2 is a Mott insulator, the symmetries of the spin density in both direct and momentum space can be captured by DFT.
-We decompose this spin density into spherical harmonics around Mn and O sites.
-In the Mn site, besides the $s$-wave staggered magnetic dipoles (see top panel in Fig 1e) we identify two ferroically ordered contributions with the symmetries of the $d_{x^{2}-y^{2}}$ and $g_{2}\equiv Y_{4,2}$ harmonics.
-Both partial waves have the same $d$-wave nodal structure, and they deform the otherwise spherically symmetric spin density.
-In the Oxygen site, which has no magnetic dipole moment, we see that its atomic spin density is fully described by a $d_{x^{2}-y^{2}}$ form factor.
-In the bottom panel of Fig.
-1e we show the $d$-wave direct-space order parameter that is obtained by adding up the $d_{x^{2}-y^{2}}$ form factors on MnA, MnB and O sublattices.
-Indeed, the spin density in the material follows the spin point group ${}^{2}4/^{1}m^{1}m^{2}m$, which classifies it as a $d$-wave altermagnet.
-The electronic band structure is spin-split (see Fig.
-2a), with spin splitting reaching a 1 eV scale, and the spin polarization of the energy iso-surfaces in momentum space (Fig.
+In Fig.
+1b we show a top view of a Lieb lattice layer in V2Te2O.
+We plot as well an iso-surface of the spin density obtained by DFT without SOC.
+This spin density follows the spin point group $^{2}4/^{1}m^{1}m^{2}m$, which classifies it as a $d$-wave altermagnet.
+Sites $A$ and $B$ are occupied by V atoms with in-plane antiparallel magnetic dipole moments.
+The radial profile of this antiferroic $s$-wave spin density is shown in the top panel of Fig 1e.
+At sites $A$ and $B$ we identify also two higher order partial waves, that are ferroically aligned (same sign in both sublattices).
+These ferroic contributions are characterized by real spherical harmonics $d_{x^{2}-y^{2}}$ and $g_{2}\equiv Y_{4,2}$, both sharing the same $d$-wave structure in the $xy$ plane, with nodes along the diagonals.
+Notably, at the $O$ site, which has no magnetic dipole moment, we see that the atomic spin density is fully described by a $d_{x^{2}-y^{2}}$ form factor.
+On the bottom panel of Figure 1e we show the radial profile of the order parameters that are constructed by adding the contributions of the $d_{x^{2}-y^{2}}$ ($g_{2}\equiv Y_{4,2}$) partial wave on $A$, $B$ and $O$ sublattices.
+Note that for this material, the ferroic order parameters are comparable in size with the $s$-wave antiferroic Néel order.
+In the Supplementary Material, we show the relative size of this atomic spin density form factors, with respect to the $s$-wave magnetic dipole, for several altermagnetic candidates with a Lieb lattice structure.
+We now focus in the electronic band structure of KV2Se2Jiang et al.
+(2025), which has the same spin density structure described before, but in a bulk tetragonal crystal.
+To understand the spin polarization of the band structure, the following minimal model is instructive Fernandes et al.
+(2024):
+$\begin{split}h(\mathbf{k})=&\lambda_{1}(k_{x}^{2}-k_{y}^{2})\sigma_{z}+\\ &\lambda_{2}(k_{y}-k_{x})k_{z}\sigma_{[110]}+\lambda_{2}(k_{x}+k_{y})k_{z}\sigma_{[\bar{1}10]}\;,\end{split}$ (1)
+where $\mathbf{k}$ is the wave-vector and $\sigma_{z}$, $\sigma_{[110]}$, and $\sigma_{[\bar{1}10]}$ are defined as $\bm{\sigma}\cdot\mathbf{v}$, with $\mathbf{v}$ a unit vector along $[001]$, $[110]$ and $\bar{1}10]$ respectively.
+$\lambda_{1}$ captures the dominant nonrelativistic $d$-wave spin splitting, and $\lambda_{2}$ ($\lambda_{2}<<\lambda_{1}$) captures additional splittings that appear when SOC is included.
+We first focus on the nonrelativistic limit ($\lambda_{2}=0$ in the previous model).
+The band structure calculated by DFT is shown in Fig.
+2a, with spin splitting reaching a 1 eV scale.
+The spin polarization of the energy iso-surfaces in momentum space (Fig.
 2c) shows the same $d$-wave character as the ferroically ordered $d_{x^{2}-y^{2}}$ projection of the spin density in direct space (Fig.
 2b).
-A key factor behind the large splitting in the A2O3M2Se2 family is the Lieb lattice structure.
-Here, the anisotropy of the exchange enters at the second Mn-Mn neighbor.
-If we consider the MnA sublattice, the links with neighbors at $(a,0,0)$ and $(0,a,0)$ are different, because the first one contains the O-site in the middle.
+A key factor behind the large splitting in this family of materials is the Lieb lattice structure.
+Here, the anisotropy of the exchange enters at the second AAA-A neighbor.
+If we consider the $A$ sublattice, the links with neighbors at $(a,0,0)$ and $(0,a,0)$ are different, because the first one contains the O-site in the middle.
 This anisotropic exchange at close distance can lead to large spin splitting and strength of the altermagnetic order.
-We emphasize that, in the non-relativistic collinear regime, both the direct-space spin density and the momentum-space spin polarization of the band structure have spin degenerate nodal planes orthogonal to the directions $[110]$ and $[\bar{1}10]$.
-When SOC is included in the DFT calculations (Néel order along $[001]$), the nodal planes show a weak $d$-wave spin-splitting, with spin polarization normal to the respective plane.
+When SOC is included in the DFT calculations (Néel order along $[001]$), the previously degenerate nodal planes show now a weak $d$-wave spin-splitting, with spin polarization normal to the respective plane (see terms proportional to $\lambda_{2}$ in Eq.
+1).
 Thus, when looking at the plane in momentum space normal to the $[110]$ axis that crosses the $\Gamma$ point (Fig.
-2i), the component $S_{110}$ shows a $d$-wave character, with nodal planes normal to the $[\bar{1}10]$ and $[001]$ axes.
+2i), the component $S_{110}$ shows a $d$-wave character, with nodal planes normal to the $\bar{1}10]$ and $[001]$ axes.
 The corresponding $S_{110}$ component of the direct-space spin density displays the same $d$-wave nodal structure (Fig.
 2f).
-Analogously, if we focus on the plane normal to the $[\bar{1}10]$ axis, both the direct-space spin density (Fig.
-2g), and the spin polarization of the energy iso-surface (Fig.
-2j) have nodal planes normal to the $[110]$ and $[001]$ axes.
-Thus, there is a clear correspondence between the nodal structure of the direct-space spin density, and the spin polarization of the band structure in momentum space.
-More broadly, as we show below, each spin component displays a different d-wave form-factor, consistent with the analysis of Ref.
-Fernandes et al.
+Analogous correspondence can be seen for the spin component $S_{\bar{1}10}$ in Fig.
+2g and Fig.
+2j.
+Even with SOC, there is a clear correspondence between the nodal structure of the ferroically ordered spin density in direct space, and the spin polarization of the band structure in momentum space.
+Each spin component displaying a different $d$-wave form-factor, is consistent with the analysis of Ref.
+Fernandes et al., 2024.
+An important consequence of the existence of two nodal planes for each of the three independent spin directions is that the band structure will only be spin-degenerate along nodal lines in the presence of SOC rather than nodal planes.
+Pure atomic altermagnet candidate—We consider Ba2CaOsO6, a double perovskite with face-centered cubic unit cell (space group $Fm\bar{3}m$) with Os at the $4a$ Wyckoff site Thompson et al.
 (2014).
-A full characterization of the partial-wave decomposition of the direct space spin density for La2O3Mn2Se2 can be seen in the SI.
-In La2O3Mn2Se2, the $[110]$ ($[\bar{1}10]$) component of the spin density is two orders of magnitude smaller than the $S_{001}$ component.
-Additional insight into the spin ordering in the presence of SOC is provided by the following minimal model:
-$\begin{split}h(\mathbf{k})=&\lambda_{1}(k_{x}^{2}-k_{y}^{2})\sigma_{z}+\\ &\lambda_{2}(k_{y}-k_{x})k_{z}\sigma_{[110]}+\lambda_{2}(k_{x}+k_{y})k_{z}\sigma_{[\bar{1}10]}\;,\end{split}$ (1)
-where $\mathbf{k}$ is the wave-vector and $\sigma_{z}$, $\sigma_{[110]}$, and $\sigma_{[\bar{1}10]}$ are defined as $\bm{\sigma}\cdot\mathbf{v}$, with $\mathbf{v}$ a unit vector along $[001]$, $[110]$ and $[\bar{1}10]$ respectively.
-Without SOC, $\lambda_{2}=0$, and when SOC is included the additional $d$-wave splitting is captured by $\lambda_{2}$, with $\lambda_{2}<<\lambda_{1}$.
-An important consequence of the existence of two nodal planes for each of the three independent spin directions is that the band structure will only be spin-degenerate along nodal lines in the presence of SOC rather than nodal planes.
-Pure atomic altermagnet candidate – We consider now Ba2CaOsO6, whose crystal structure is face-centered cubic with space group $Fm\bar{3}m$ (No.
-225) Thompson et al.
-(2014).
-The Os atom occupies the Wyckoff position $4a$.
-This material has drawn significant attention owing to unexpected oscillations in the zero-field muon spin relaxation Thompson et al.
+This material attracted significant attention due to unexpected oscillations in zero-field muon spin relaxation Thompson et al.
 (2014), which suggest a phase transition with TRSB at $T\sim 50\;K$.
@@
 Later on, this claim was supported from an ab initio perspective Fiore Mosca et al.
-(2022), using VASP and on-site density matrix control Allen and Watson (2014).
-To our knowledge, however, little has been said about the connection between this ferro-octupolar phase and the spin splitting in the electronic band structure.
-Despite SOC being strong in this double perovskite, it is still instructive to explore its non-relativistic limit, because also in this limit we find a correspondence between the direct and momentum space order parameters.
-We calculate the spin density from DFT, including correlations on the Os site using the Liechtenstein’s spherically symmetric DFT+U approach Liechtenstein et al.
-(1995), with parameters $U=3.2$ eV and $J=0.5$ eV, in agreement with the pioneer study on the compound Fiore Mosca et al.
 (2022).
+To our knowledge, the connection between this octupolar state and the spin splitting in the band structure remains unexplored.
+Despite strong SOC in Ba2CaOsO6, the non-relativistic limit remains instructive, as we still find a correspondence between direct- and momentum-space order parameters.
+We compute the spin density using DFT+U (Liechtenstein scheme) with $U=3.2$ eV and $J=0.5$ eV, following the approach of Ref.
+Fiore Mosca et al., 2022.
 Without SOC, the collinear spin density around the Os site is shown in Fig.
 1c.
-Below, in Fig.
-1f we see that the harmonic decomposition is given by $d_{xy}$ and $g_{\bar{2}}\equiv Y_{4,-2}$.
-Both spherical harmonics have $d$-wave symmetry in the $xy$ plane, but unlike La2O3Mn2Se2, the nodal planes are normal to $[100]$ and $[010]$.
-There is no magnetic dipole moment, as can be seen in the top panel of Fig.
-1f, making this material a pure atomic altermagnet.
-The corresponding iso-energy surfaces (Fig.
-3c) follow the $[C_{2}||C_{4z}]$ spin symmetry.
+Its partial wave expansion (Fig.1f) reveals $d_{xy}$ and $g_{\bar{2}}\equiv Y_{4,-2}$ components, with nodal planes along $[100]$ and $[010]$.
+No magnetic dipole moment is present, identifying Ba2CaOsO6 as a pure atomic altermagnet, with spin point group $^{2}4/{}^{1}m{}^{2}m{}^{1}m$.
+The iso-energy surfaces (Fig.3c) exhibit the same nodal structure as the spin density in direct space.
 In Fig.
-3a we show the band structure along a path in the $(k_{x},k_{y})$-plane that bisects the two nodal planes.
+3a we show the band structure along a path in the $k_{z}=0$ plane that bisects the two nodal planes.
 We identify a spin splitting of 0.2 eV just below the Fermi level.
-We emphasize that, because Ba2CaOsO6 is a Mott insulator, the bands crossing the Fermi level in Fig.
+We emphasize that, because Ba2CaOsO6 is a Mott insulator, bands crossing the Fermi level in Fig.
 3a will be pushed away by the on-site repulsion.
-Further investigation of the impact of strong correlations on the band structure is, however, beyond the scope of our work, which focuses on the symmetry-enforced shape of the spin density on the Os site.
+A full treatment of strong correlations is beyond the scope of this work, which focuses on the symmetry-enforced shape of the spin density.
 In this regard, we note that the $d_{xy}$ form factor of the spin density obtained via DFT+U is consistent with the strong-coupling perspective on the octupolar moment of Os $5d^{2}$, which can be understood as the $z$-component of the pseudo-spin that characterizes the non-Krammers doublet ground-state of the Os atom, arising from the combined effects of crystal fields and SOC on the localized $J=2$ state Paramekanti et al.
@@
 (2021).
-Figure 3: DFT+U calculation for Ba2CaOsO6 without (a-c) and with (d-j) SOC .
-(a), (d) Band structure colored according to the $S_{z}$ spin polarization.
+Figure 3: DFT+U results for Ba2CaOsO6 without (a–c) and with (d–j) SOC.
+(a,d) Band structures colored by $S_{z}$ spin polarization.
 (b) Iso-surface of the direct-space spin density (no SOC) around the Os site.
-(c) Energy iso-surface cut in the plane $k_{z}=0$.
-Iso-surface of the spin density around the Osmium site (with SOC), projected along the spin components $[100]$ (e), $[010]$ (f) and $[001]$(g).
-Energy iso-surface cuts in planes that contain the $\Gamma$ point and are normal to directions $[100]$ (h), $[010]$ (i) and $[001]$(j), with color scale representing the spin polarization along the axis normal to the planes.
-Once SOC is included, the spin density in Ba2CaOsO6, in agreement with previous reports Fiore Mosca et al.
-(2022), becomes strongly noncoplanar, with equivalent $d$-wave form factors along the $x$, $y$, and $z$ axes.
-These results are consistent with the analysis of Ref.
-Fernandes et al.
-(2024), which found that the inclusion of SOC makes the spin-density noncoplanar in some altermagnets.
-Figure 3d shows the band structure (with SOC) along a path in the $x-y$ plane, bisecting the planes normal to $[100]$ and $[010]$ axes.
-We also show the $z$-component of the spin polarization in a color scale.
-Note that, in the presence of SOC, the system displays a band gap of around 1 eV.
-We note, however, that strong correlations will render the material a Mott insulator.
-The $x$, $y$ and $z$ components of the spin density around the Os atom are shown in Figs.
-3 e-g.
-Below each spin density, in Figs.
-3 h-j, we show energy iso-surface cuts with the spin polarization (component normal to each plane) plotted in color scale.
-The spin density can be understood as 3 orthogonal $d$-wave form factors corresponding to the $x$, $y$ and $z$ spin components.
+(c) Energy iso-surface in the $k_{z}=0$ plane.
+(e–g) Spin density iso-surfaces with SOC, projected along $[100]$, $[010]$, and $[001]$.
+(h–j) Energy iso-surface cuts in planes normal to $[100]$, $[010]$, and $[001]$, colored by spin polarization normal to each plane.
+Upon inclusion of SOC, the spin density in Ba2CaOsO6 becomes strongly noncoplanar, with equivalent $d$-wave form factors along $x$, $y$, and $z$, consistent with Refs.
+Fiore Mosca et al., 2022 and Fernandes et al., 2024.
+Figure 3d shows the band structure (with SOC) along a path in the xyx-y plane, bisecting the planes normal to $[100]$ and $[010]$ axes.
+The $z$-component of the spin polarization is shown in color scale.
+SOC opens a  1 eV band gap.
+We note, however, that strong correlations will make the system a Mott insulator.
+In analogy with Eq. 1, a simple model capturing these noncoplanar altermagnetic $d$-wave form factors is given by Fernandes et al. (2024):
+[…]
+Note that, because of the cubic symmetry of the crystal structure, and in contrast to KV2Se2O, all spin components have the same prefactor $\lambda$.
+The DFT spin density projected along $x$, $y$, and $z$ is shown in Figs.
+3e–g.
+Below, Figs.
+3h–j show energy iso-surfaces with spin polarization (component normal to the plane) in color.
+Comparing the DFT calculations with Eq.
+2, we can understand the spin density as 3 orthogonal $d$-wave form factors corresponding to the $x$, $y$ and $z$ spin components.
 The nodal structure of the spin density is reflected in the spin polarization of the energy iso-surfaces.
-In analogy with Eq.
-1, a simple model capturing these noncoplanar altermagnetic $d$-wave form factors is given by Fernandes et al.
-(2024):
-$h(\mathbf{k})=\lambda k_{x}k_{y}\sigma_{z}+\lambda k_{z}k_{x}\sigma_{y}+\lambda k_{y}k_{z}\sigma_{x}\text{ .}$ (2)
-Note that, because of the cubic symmetry of the crystal structure, and in contrast to La2O3Mn2Se2, all spin components have the same prefactor $\lambda$.
-The emergence of the atomic $d$-wave form factors in Ba2CaOsO6, with or without SOC, comes from an interplay between correlations and orbital ordering.
+The $d$-wave form factors in Ba2CaOsO6 arise from an interplay between correlations and orbital order.
 The on-site occupation matrix of the $d$ orbitals captures the $d$-wave character Fiore Mosca et al.
 (2022), which then gets imprinted in both the direct-space spin density and the spin splitting in the band structure.
-Figure 4: Estimation of the spin-splitter angle for the non-relativistic calculation of both La2O3Mn2Se2 (a) and Ba2CaOsO6 (b).
-See in Supplementary Information the calculation of the GMR coefficients.
-Spin currents – We calculate the magnitude of the spin currents in the non-relativistic limit for both La2O3Mn2Se2 and Ba2CaOsO6, by quantifying the spin-splitter angle $\alpha$, which measures the efficiency in the charge-to-spin conversion, and the GMR coefficient Šmejkal et al.
+A similar orbital-order mechanism was recently studied in a Hubbard square lattice model Giuli et al.
+(2025).
+Figure 4: Spin-splitter angle from non-relativistic DFT for KV2Se2O (a) and Ba2CaOsO6 (b).
+Spin currents – We compute spin currents in the non-relativistic limit for KV2Se2O and Ba2CaOsO6.
+We quantify the spin-splitter angle $\alpha$ González-Hernández et al.
+(2021) and the GMR coefficient Šmejkal et al.
 (2022c).
-Our results are shown in Fig.
-4, where the spin-splitter angle is plotted as a function of the Fermi energy.
-In the explored energy interval, the spin-splitter angle for Ba2CaOsO6 (Fig.
-4a) reaches a maximum of 26 , while for La2O3Mn2Se2 it is even higher, reaching 44.
-The GMR coefficient (see SI) in La2O3Mn2Se2 reaches 900% for energies close to the Fermi energy, which is one order of magnitude higher than the 100% scale of the initial altermagnetic GMR prediction, considering RuO2 Šmejkal et al.
+Figure 4 shows $\alpha$ as a function of Fermi energy, peaking at 26 for Ba2CaOsO6 and 42 for KV2Se2O.
+The GMR coefficient in KV2Se2O reaches 800% (see SI), nearly an order of magnitude above original prediction in RuO2 Šmejkal et al.
 (2022c).
-While La2O3Mn2Se2 is not metallic, we expect the spin-splitter angle and the GMR coefficient to be equally large in related metallic altermagnetic candidates on layered Lieb lattices, such as UCr2Si2Lemoine et al.
-(2018), KV2Se2O Jiang et al.
-or RbV2Se2O Zhang et al.
+Comparable values are expected in related metallic Lieb-lattice altermagnets such as RbV2Se2O Zhang et al.
 (2025).
-In addition to the $d$-wave form factor in the sites hosting the magnetic dipoles, the $d$-wave atomic spin density will be present, without dipole, at the O site in RbV2Se2O and KV2Se2O, and at the U site at UCr2Si2C.
-In the SI, we show calculations of the electronic structure for UCr2Si2C, and we comment as well on the nature of the spin conductivity tensor in Ba2CaOsO6(with SOC), when the spin density is described by three noncoplanar $d$-wave form factors.
-Discussion – While the notion of multipolar magnetism is naturally connected with altermagnetism Bhowal and Spaldin (2024); Fernandes et al.
+In the SI, we comment on the nature of the spin conductivity tensor in Ba2CaOsO6  when the spin density is described by three noncoplanar $d$-wave form factors.
+Discussion—While the notion of multipolar magnetism is naturally connected with altermagnetism Bhowal and Spaldin (2024); Fernandes et al.
 (2024), one should be careful when comparing these two concepts.
@@
 The interplay between strong correlations and orbital occupation (possibly with strong SOC) generates a spin density which lacks magnetic dipole moments, while having a $d$-wave component which is also reflected in the corresponding spin-polarization symmetry of the electronic structure in the momentum space.
-In La2O3Mn2Se2 we demonstrate the coexistence of dipole and higher-order partial-wave components on the Mn-sites, with the absence of the dipole and the presence of the higher-order partial-wave component (atomic altermagnetism) on the O-sites.
-We estimate the spin-splitter angle and the GMR coefficient in this material as 44 degrees and 900%, respectively.
 We identified pure atomic altermagnetism with $d$-wave symmetry in the non-relativistic limit of Ba2CaOsO6.
-When including SOC, the spin density has the structure of three noncoplanar (and orthogonal) $d$-waves form factors.
+When including SOC, the spin density has the structure of three noncoplanar (and orthogonal) $d$-wave form factors.
 In all cases, we identify one-to-one correspondence between the ferroically ordered $d$-wave harmonic of the spin density in direct space, and the angular dependence of the spin splitting around the $\Gamma$ point.
-[…]
+In a family of compounds in the Lieb lattice (e.g.
+KV2Se2O) we demonstrate the coexistence of dipole and higher-order partial-wave components on the V-sites, with the absence of the dipole and the presence of the higher-order partial-wave component (atomic altermagnetism) on the O-sites.
+We estimate the spin-splitter angle and the GMR coefficient in this material as 42 degrees and 800%, respectively.
+This work thus extends the search for altermagnetic candidates to materials that, while breaking time-reversal symmetry, do not have dipolar order.
+The detection of the magnetic order without local magnetic dipoles might represent an intriguing challenge for the conventional techniques like neutron scattering.
+The spin-polarized electronic structure, and the spintronic responses associated with the altermagnetic order, can serve as experimental probes of these higher-order partial-wave forms of magnetic ordering.
+Besides Ba2CaOsO6, there is a range of materials that have been previously identified as candidates for magnetic octupolar order that may deserve further investigation, in the context of the higher-partial-wave ordering and atomic altermagnetism Santini and Amoretti (2000); Lovesey et al.
+(2003); Aoki et al.
+(2008); Matsumura et al.
+(2009); Sibille et al.
+(2020), including the metallic cage compound PrV2Al20, which has recently drawn significant attention Tsujimoto et al.
+(2014); Freyer et al.
+(2018); Patri et al.
+(2019); Sorensen and Fisher (2021); Ye et al.
+(2024).
 Acknowledgments—RJU, VKB, RZ, JS and LS acknowledge funding by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation)- TRR 173– 268565370 (project A03) and TRR 288– 422213477 (project A09 and B05).
@@
 FA9559-21-1-0423.
+The authors acknowledge the computing time granted on the supercomputer Mogon at Johannes Gutenberg University Mainz (hpc.uni-mainz.de).